A 2.144-g sample of phosgene, a compound used as a chemical welfare agent during World War 1, Contains 0.260g of carbon, 0.347g of oxygen, and 1.537g of chlorine. what is the empirical formula of this substance



1
Expert's answer
2022-05-24T04:18:07-0400

CxOyClz

n(C) = m(C)/Mr(C) = 0.260 g / 12 g/mol = 0.02 mol

n(O) = m(O) /Mr(O) = 0.347 g / 16 g/mol = 0.02 mol

n(Cl) = m(Cl) / Mr(Cl) = 1.537g / 35.5 g/mol = 0.04 mol

n(C) : n(O) : n(Cl) = 0.02 : 0.02 :0.04 = 1 : 1 : 2


empirical formula COCl2